### 30 Jan 2024 v2

# Foregrounding Artist Opinions: A Survey Study on Transparency, Ownership, and Fairness in AI Generative Art

Abstract.
Generative Artificial Intelligence (AI) tools are used to create art-like outputs and aid in the creative process. While these tools have potential benefits for artists, they also have the potential to harm the art workforce and infringe upon artistic and intellectual property rights. Without explicit consent from artists, Generative AI creators scrape artists’ digital work to train Generative AI models and produce art-like model outputs at scale. These outputs are now being used to compete with human artists in the marketplace as well as being used by some artists in their generative processes to create art. We surveyed 459 artists to investigate the tension between artists’ opinions on Generative AI art’s potential utility and harm. This study surveys artists’ opinions on the utility and threat of Generative AI art models, fair practices in the disclosure of artistic works in AI art training models, ownership and rights of AI art derivatives, and fair compensation. We find that artists, by and large, think that model creators should be required to disclose in detail what art and images they use to train their AI models. We also find that artists’ opinions vary by professional status and practice, demographics, whether they have purchased art, and familiarity with and use of Generative AI. We hope the results of this work will further more meaningful collaboration and alignment between the art community and Generative AI researchers and developers.
AI Art, Generative AI, AI Ethics, Survey Study, Consent, Privacy
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 3-6, 2024; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/18/06
Figure 1. “Art is the joint responsibility of all [people]”  (Huyghe, 1970).
An illustration of a rose garden connecting the minds of artists and being pruned by a machine
Figure 1. “Art is the joint responsibility of all [people]”  (Huyghe, 1970).
1. Introduction
Art has a rich history in co-evolution with the development of new technologies (Etro and Galenson, 2023). The advent of Artificial Intelligence (AI) tools used to create or assist in the creative process carries great potential utility for artists but also harbors the significant potential for harm to the art workforce and artistic and intellectual property. To examine this tension between utility and threat, we surveyed 459 artists about their views on AI-generated art.
Our study surveys artists’ opinions on fair practices in the disclosure of artistic works in AI training models, ownership and rights of AI art derivatives, and fair compensation models. Future work will involve conducting long-form interviews with them to understand artists lived experiences. Our research questions for this study can be summarized as follows: RQ1: Do artists see AI models as a threat to art workers?; RQ2: Do artists see AI art models as a positive development in the field of art?; RQ3: What types of disclosure for the use of artwork to train an AI art model do artists consider fair?; RQ4: From an artist’s perspective, who should own the derivative artwork created by AI art models?
We summarize our research results as follows: (Q1) 61.87% of survey participants agree that AI models are a threat to art workers. (Q2) 44.88% of survey participants agreed that AI art models are a positive development in the field of art (22.03% of survey participants agreed that both AI was a positive development and a threat). (Q3) Artists overwhelmingly agreed (80.17%) that model creators should be required to disclose in detail what art and images they use to train their AI models. (Q4) Artists were relatively split between thinking that derivatives from AI art models should be owned by the original artists whose work is represented in the AI art (41.39%) and the user who used the model to create the AI art (39.22%). In our study, 26.80% of the participants agreed that the work and its derivatives be considered the property of the AI model creators. We also looked into what types of compensation artists consider to be fair in exchange for the use of their artwork to train an AI art model. Half of the respondents (50.97%) stated that they did not need compensation, but what mattered to them was who was profiting from the use of their artwork. The most common fair compensation option selected by the survey participants was “I don’t need to profit, but I don’t want for-profit companies to profit from my art” (22.80%). This suggests that artists are not entirely against their art being used in AI art model training data but oppose for-profit companies benefiting financially. We describe the survey results in detail in Section 3 and determine which characteristic variables of our survey participants are most associated with these results. We will present the survey methodology in Section A.1.
Our paper draws on our team’s multidisciplinary background – in art and art history, social science, complex systems, computer science, mathematics, and statistics – to address these questions and understand the possible impact of Generative AI on artists. We hope this work will provide a modest step towards understanding artists’ opinions on Generative AI technology, its impact on their field, and the role they would like their work to play in the burgeoning field of Generative AI.
2. Background
There is ongoing disagreement regarding whether the output of Generative AI can be considered art (AI, 2023a; Commission, 2023; Jiang et al., 2023); we will not examine the topic in depth here. For the sake of convenience, we will refer to the comparable output of AI as “art” throughout this paper. However, this is not intended to refute (or endorse) the claim that AI cannot create art.
According to work by Jiang et al. (Jiang et al., 2023), “following philosophers of art and aesthetics from varied disciplines… we define art as a uniquely human endeavor connected specifically to human culture and experience.” They distinguish between inspiration/training and technique/creation. They also point out that a person being able to mimic another person’s artistic style so perfectly as to be indistinguishable is rare, as good forgeries are typically noteworthy. They claim that these aspects, among others, differentiate between what humans and AI can create.
Another dimension of debate concerning whether Generative AI creates art is whether the output of the model is copyrightable under US law. The Courts in the United States recently rejected for copyright a work the plaintiff described as “autonomously created by a computer algorithm running on a machine,” saying there is a “consistent understanding that human creativity is the sine qua non at the core of copyrightability, even as that human creativity is channeled through new tools or into new media” (US District Court for the District of Columbia, 2023).
Among the arguments intended to distinguish human art from AI-generated art, we find three that are particularly compelling. First, human art is the product of activity without “organic pressure” – allowing us to distinguish between the music of people, songbirds, and AI models (Jiang et al., 2023). Second, that the world and the artist are always tightly coupled (Jiang et al., 2023), in part through corporeality (Pfeifer and Bongard, 2006). Third, that art plays a special role in the lived experience of humans (Huyghe, 1970).
Regardless of whether AI can make art, the output of Generative AI can meet the minimally viable criteria for a commercial product, like the cover of a book (Crabapple, 2022) (see Sec. 2.3), which means it is making its way into some of the many places art currently lives in our world, impacting artists and others along the way.
It is important to note that, although how technology can be incorporated into art may be up for debate, that technology is part of the artistic process is not new or ambiguous. Artists have always used tools (Huyghe, 1970), and many art forms involve machines (such as woodwork, metalwork, and digital art) and automation (such as digital printers and cameras). Some artists also choose to use technology in ways other than its intended purpose or develop new art-making techniques by breaking technology. Such artists may be both suspicious and intrigued by the technology they use, questioning the relationship between humans and machines and expanding art practices and processes. When presented with new technology, an artist may ask questions like “How can I keep my work with the machine intentional? How might this technology distract me? Is the machine using me, or am I using it? Is this image-making or art-making? How does this automation serve my artistic intentions and practice?” (Karson, 2023).
In the case of machine learning, a crucial question for many artists is “What data was used to train this model?” (See Sec. 3.) Such inquiries are part of the artistic tradition of challenging the status quo and are in stark contrast to how the creators of Generative AI models have presented their tools for use to the general public (AI, 2023a).
2.1. Generative AI
The field of Generative AI has created much excitement around its relevant applications, both anticipated and emergent (Harby, 2023; Bubeck et al., 2023; Abdin et al., 2023). Its valuation of tens of billions of dollars (Jiang et al., 2023; for Human-Centered Artificial Intelligence, 2023; Crabapple, 2022) has, of course, inspired a gold rush of research (academic, industrial, and otherwise) and media coverage (Shan et al., 2023; Gozalo-Brizuela and Garrido-Merchán, 2023). Many thousands of models are available and accessible (Face, 2023), along with an ecosystem of tools supporting their use. (Gozalo-Brizuela and Garrido-Merchán, 2023) The space of related work being explored, therefore, extends from well-established benchmarking (Abdin et al., 2023; Lynch, 2023; for Human-Centered Artificial Intelligence, 2023), all the way to cynical (Nicoletti and Bass, 2023), often predatory (Hall, 2023), money-making schemes and fraud (Horton, 2023). Since 2010, the number of yearly AI-related publications has only grown, reaching nearly 500,000 in 2021 alone (for Human-Centered Artificial Intelligence, 2023).
The core functionality of a Generative AI model is as follows: based on input from a person – through a user interface, as a conversational text prompt, or as a piece of media to use as a starting point, such as a video, text, code snippet, or an image – Generative AI models can produce novel output in an equal variety of media (see Fig. 2(Gozalo-Brizuela and Garrido-Merchán, 2023).
Currently, the major players in textual Generative AI are large language models (LLMs), such as OpenAI’s ChatGPT and GPT-4 (Zhao et al., 2023; AI, 2023b). These models work by learning how to provide a plausible next word in response to a sequence of text, which results in seemingly fluent linguistic output (Zimmerman et al., 2023). In the visual sphere, models like StableDiffusion and Midjourney (Holz and team, 2024) can modify an image by iteratively tweaking random noise until they create an image that matches the user’s input. The model decides which part of the image needs to be modified at any given time based on the diffusion process (Andrew, 2023).
For the purposes of our paper, we define Generative AI as being a subset of tools or processes that combine massive probability distributions with sense-making algorithms in order to create output that, taken as a whole, was not included in the training data (Jiang et al., 2023). The probability distribution should be based on the training data. The sense-making algorithm could be based on several strategies, such as a physical process like diffusion or an architectural approach like transformers/neural nets, although in the end, it should produce something akin to a latent/embedding space (Shan et al., 2023); a place to hold meaning (Gozalo-Brizuela and Garrido-Merchán, 2023; Zhao et al., 2023; Grand et al., 2022). Research is ongoing into the various distributions and processes that can be implemented in Generative AI models to achieve specific behaviors and results (Xu et al., 2023). However, irrespective of details, all of these models use advanced architectures in conjunction with large amounts of data gathered through brute-force methods across vast swathes of the publicly accessible internet  (Xu et al., 2023; Zhao et al., 2023; Salkowitz, 2022).
Figure 2. Generative AI models can produce novel output in various media such as video, text, code snippet, or image. This output is based on the input from an individual through a user interface, conversational text prompt, or a piece of media, as mentioned in a survey conducted by Gozalo Brizuela (Gozalo-Brizuela and Garrido-Merchán, 2023). The technology is advancing rapidly, and in many contexts, the output generated by these models is indistinguishable from that of a human.
A schematic of typical inputs to Generative AI models such as text and images and the resulting outputs (e.g., text, images, video, code, music)
Figure 2. Generative AI models can produce novel output in various media such as video, text, code snippet, or image. This output is based on the input from an individual through a user interface, conversational text prompt, or a piece of media, as mentioned in a survey conducted by Gozalo Brizuela (Gozalo-Brizuela and Garrido-Merchán, 2023). The technology is advancing rapidly, and in many contexts, the output generated by these models is indistinguishable from that of a human.
Here, we are interested specifically in how artists perceive Generative AI, what its impact on art-related careers might be, and what moral, ethical, artistic, and financial implications are coming hand-in-hand with the broader flourishing of Generative AI (Shan et al., 2023; Baio, 2022; Crabapple, 2022; Murphy, 2022; for Human-Centered Artificial Intelligence, 2023; Gozalo-Brizuela and Garrido-Merchán, 2023; Zhao et al., 2023). We, therefore, touch on four relevant topics: first, related academic work; second, the economic context for this discussion; third, what sort of stages are within the pipeline of Generative AI models and how those stages can each cause harm (a specific and noteworthy type of impact); fourth, an example of how a previous technological change caused upheaval within a creative field.
Although in our paper we focus more on the pressing potential for harm, it is worth noting that there is also growing excitement around Generative AI’s capabilities and how it could be incorporated into artistic processes (Karson, 2023). For example, artist Cosmo Wenman worked with blind researchers Brandon Biggs, Lindsay Yazzolino, and Joshua Miele to create images using the Midjourney tool (Holz and team, 2024). Their team noted that programs like Midjourney can find “weird and thematic intersections” in art that go unnoticed by human eyes. They also discussed the role of chaos in creating “visual intrigue” and how technology can be used to “enhance agency.” Wenman excitedly called the new technology “essentially moveable type for visual rhetoric” (Schrader, 2023).
2.2. Related work
Jiang et al.(Jiang et al., 2023) analyze Generative AI art models and their potential impact on the world through the lens of various academic sources (both philosophical and scientific). They describe many opportunities for harm as a result of such products: violation of the consent and privacy of artists (Crabapple, 2022; Murphy, 2022; Baio, 2022; Ho, 2024; the Midjourney team, 2023), a lack of dataset transparency (shielding violations and biases) (Ho, 2024; the Midjourney team, 2023), a failure to uphold the contextual integrity of content (Nissenbaum, 2004), harm to the livelihood of the artist, biases in the training data, automating processes in a way that is detrimental to the creation of art, reduced accessibility, erasure of human works by obscurity, and increasing reluctance by artists to share their works openly (Crabapple, 2022).
The harms Jiang et al. (Jiang et al., 2023) foresee for artists fall into two categories, financial and emotional, which manifest primarily as the replacement of a human artist on a project, the impersonation of an artist without consent (Baio, 2022; Murphy, 2022; Crabapple, 2022; Gozalo-Brizuela and Garrido-Merchán, 2023), or plagiarism. Impersonation, particularly in this context, could be weaponized for nefarious purposes (Horton, 2023; Ho, 2024; the Midjourney team, 2023). Moreover, these two kinds of harm, enacted against individuals at scale, create societal-level harm. Specific societal harms Jiang et al. discuss in their work include data laundering – through academic or open-source projects where public funds are funneled into research with ultimately private profits – the proliferation of stereotypes and “hegemonic views,” and erasure by obscurity (i.e., that AI-generated outputs can be created much more quickly than human ones). They also point out that US copyright and intellectual property jurisprudence needs to be better equipped to deal with these new tools and that open sourcing, often called on as a solution to software-related problems, is incomplete in this case (Jiang et al., 2023).
As possible solutions, Jiang et al. bring up regulations regarding opt-in consent (for training data), independent funding (much current research is supported by industry), and education about how power interacts with technology. In particular, they note there are many incentives to label work as “apolitical” and that the spin capitalism is putting on Generative AI is the “democratization of art” (and other fields) (Jiang et al., 2023).
Going beyond Jiang et al.’s list of solutions, Shan et al.  (Shan et al., 2023) propose tools allowing artists to “poison” models that scrape their work without permission. The tool they created (Nightshade) can “destabilize general features” in the targeted generative model and “bleed-through” to semantically similar prompts without injecting large amounts of tailored data (Shan et al., 2023).
2.3. Economic context: livelihood of the artist
AI (a term which includes, but is not limited to, Generative AI) is increasingly working its way into the world economy. Across nearly every economic sector for which data is gathered, AI-related job postings have been increasing (for Human-Centered Artificial Intelligence, 2023). Globally, employers in a broad range of fields seek AI-skilled workers, and every year, many companies decide to adopt AI tools. The top benefit these companies claim to see in response to that decision is lowered costs (for Human-Centered Artificial Intelligence, 2023), including through layoffs and replacing employees with AI (Korn, 2024). On the other end of the employment ladder, workers in an increasing variety of fields list AI as a relevant skill (for Human-Centered Artificial Intelligence, 2023).
Some of these impacts are already coming to bear on the economic reality of artists. Entry-level positions in film, TV, and gaming seem particularly at risk, and the conceptual period of a project seems to be a popular phase in which to replace artists with AI (Carter, 2023). Traditional niches like illustrating book covers can now be filled by the output of Generative AI (Crabapple, 2022; Chalk, 2024), and major employers can reduce the number of employees needed (for example, Netflix Japan’s layoff of animators) (Jiang et al., 2023).
It isn’t just the output of AI models that can have an economic impact on artists, however. Harm, both financial and emotional, can occur based on whether people have – or even could have – consented to the content they created and shared in a certain context (Nissenbaum, 2004) (especially in the complex context of social media, e.g., an artist with a physical gallery presence whose work was photographed by a visitor) being repurposed by machine learning models and the people – and companies – who control or use them (Solon, 2019; Baio, 2022; Ho, 2024; the Midjourney team, 2023). The harm done by the AI model can be incipient before the AI model was ever conceived of; the complications of the model creation pipeline are discussed in more detail in Sec. 2.4.
Furthermore, the issue of copyright protection for artworks (made by humans) to train commercial AI models (without consent, attribution, or sufficient compensation to human authors) is a matter of concern for artists, creative industries, and federal agencies (Commission, 2023; Cho, 2023; Team, 2024; Patel and Palazzolo, 2024; David, 2024; Marcus and Southen, 2024).
2.4. Problems in the pipeline of Generative AI tools
As mentioned, when we consider the impact of an AI model, the impact is not just how the model’s output is used or how the model itself is deployed (as a replacement for human workers, etc.). Each stage within the pipeline of creating a Generative AI model can contribute to the eventual harm done by (and therefore the impact of) the model, even though most stages, with the sometimes exception of the two endpoints (input and output), are relatively inaccessible. In this section, we outline the pipeline of creating a Generative AI model, and we examine one particularly pervasive harm, bias (Zhao et al., 2023; Gallegos et al., 2023), in order to explain how multiple stages within the pipeline can contribute to the model’s impact.
When creating an AI model, training materials must be gathered (often a subset of artifacts from the real world) (Salkowitz, 2022); those materials must be groomed; poor quality, irrelevant, or over-represented artifacts should be removed or otherwise addressed (Luccioni and Viviano, 2021; Zhao et al., 2023). The architecture of the model itself must be assembled. The model must be trained. For complex Generative AI models whose behavior can be unpredictable, there is the additional process of alignment, which is intended to bring the model’s range of possible — or at least likely – behavior in line with whatever the creators of the model deem desirable (Zhao et al., 2023).
Very early in the pipeline, choosing the training material has a significant influence on the model’s eventual behavior. Important datasets used as benchmarks for machine learning have been found to be “overwhelmingly composed of lighter-skinned subjects” (Buolamwini and Gebru, 2018), introducing bias obliquely into the model (the global bias introduced would not be caught by locally vetting each piece of training data for harmful content). However, there is also explicitly objectionable and biased content in the datasets themselves – such as racist hate speech and harmfully sexualized images (Luccioni and Viviano, 2021; Cole, 2023) – even after cleaning procedures.
Midway through the pipeline, in the interior of the trained model, pieces as basic and abstracted as the meaning of common English words have been found to encode gender biases (Bolukbasi et al., 2016). Near the end of the pipeline, the output produced by the model can instantiate bias. For example, the Stable Diffusion model disproportionately generates images of white male executives, doctors, lawyers, judges, and law enforcement, and also disproportionately generates images of black and brown criminals and service workers (Nicoletti and Bass, 2023; Ghosh and Caliskan, 2023). Additionally, some users have found that Generative AI tools readily produced images in which they looked younger, more sexualized, or whiter, even in cases where their prompt did not request or hint at those parameters (Snow, 2022; Ghosh and Caliskan, 2023; Hall, 2023).
These examples are not unique; similar biases are found across current AI models and products – including those of Generative AI – at every level, from the training data they learned from to their internal representations to their outputs (Luccioni and Viviano, 2021; Ghosh and Caliskan, 2023; Cheng et al., 2023; Gallegos et al., 2023; Zhao et al., 2023; Common Crawl, [n. d.]; Cole, 2023).
As a final problem in the pipeline, it is worth noting that these AI tools cannot be created ex nihilo: they are integrated into complex human systems that bring their own forces to bear on the model’s potential impact. There are discrepancies in the demographics of who is wielding the power of these new technologies: who is making them (for Human-Centered Artificial Intelligence, 2023), profiting off them, using them. This can have many knock-on effects, sometimes through mechanisms difficult to foresee (Even-Tov et al., 2023), with respect to the design, use, and role of these tools in society (Jiang et al., 2023; for Human-Centered Artificial Intelligence, 2023). In particular, the companies behind these products often employ a great deal of secrecy as to their training data, architecture, or implementation (or all three) (Ho, 2024; the Midjourney team, 2023). As a result, users must interact with the product as a black box (Jiang et al., 2023; Marcus and Southen, 2024); concerned users can only conjecture as to what kind of training data could produce the observed results based on what is known of the basic architecture of the model, which in turn allows argument over the obscure details (Team, 2024; Masse, 2024). The lack of transparency benefits the companies offering these products (with respect to establishing the impact, specifically the harms, of the model) in giving them a way to sidestep moral and ethical concerns (Marcus and Southen, 2024; Masse, 2024).
2.5. Previous disruption: Napster’s example
Similarities exist between the early years of the Peer-to-Peer music sharing platform Napster and the nascent stage of Generative AI tools. Both technologies were exciting and disruptive, with the potential for financially high stakes (Jiang et al., 2023; for Human-Centered Artificial Intelligence, 2023; Bubeck et al., 2023; Bergen, 2002). Both also forged ahead despite glaring questions of legality, particularly with respect to copyright (Grynbaum and Mac, 2023). In Napster’s wake, legal scholars debated whether contemporaneous copyright law did or should apply to the digital age, and they wondered whether the public’s perception of copyright laws had changed due to Napster’s large user base and the demand for free or reduced-cost music (Ku, 2002; Bergen, 2002). Ultimately, Napster shut down, having underestimated the legal battles it would face (Green, 2002; YellowBrick, 2023).
The argument that Generative AI technologies should be able to be “taught” their downstream abilities (to speak a language, to reason, to create various output) (Masse, 2024; Team, 2024; team, 2023) by much of the (ostensibly (Marcus and Southen, 2024)) publicly available Internet (Salkowitz, 2022) – or that, morality aside, to prevent them would be infeasible – is a more complex version of the argument made by Napster’s founders and early users who insisted that music should (or would) be available to the public for free (Marcus and Southen, 2024).
In the face of such arguments, it is worth noting that openness is not a good in and of itself, and can lead to its own harms, such as violations of contextual integrity or privacy (Nissenbaum, 2004). Furthermore, the forces, often economic, that shape underlying practices like web scraping (Common Crawl, [n. d.]; Salkowitz, 2022) may not be the forces best suited to beneficently shaping technological development (AI, 2023b).
The apparent success of music distribution platforms such as Apple Music and Spotify (Wise, 2023) – and the absence of Napster – suggests that there is significant doubt as to the eventual triumph of Generative AI models over pending legal challenges. As of now, at least, there are no exceptions for AI companies regarding US copyright law (Commission, 2023).
3. Results
Our anonymous survey of artists is an observational study (N=459) of artists over the age of 18 years (we report participant demographics in Table 1). We see from the responses to our primary research questions (see Table 2) that 80.17% of the survey participants agree that model creators should be required to disclose in detail what art and images they use to train their AI models. We also find that 61.87% of survey participants agree that AI models are a threat to art workers. In contrast, 44.88% of survey participants agree that AI art models are a positive development in the field of art, notably with 27.02% of artists selecting neutral on the Likert scale for this question. Interestingly, we see an overlap between artists who agree that AI art is a positive development and those who also agree that it is a threat to art workers; 22.03% of survey participants agreed that both AI was a positive development and a threat.
We look at the question that asks participants to think about a scenario where an AI art model was used by someone else to produce artwork recognizably in their (the survey participant’s) style and think about who should own the work and its derivatives. Only 26.80% of the participants agreed that the work and its derivatives be considered the property of the AI model creators, while 39.22% of participants agreed that the work and its derivatives be considered the property of the person who used the AI model to generate the artwork. Finally, 41.39% of the participants agreed that the work and its derivatives be considered the property of the artists whose style is being represented by the AI model output. In the following sections, we will analyze which variables are most associated with participants in our survey study who agreed with the primary research questions.
Type Count (%)
Male 273 (59.48%)
Female 163 (35.51%)
Non-binary 23 (5.01%)
White 361 (78.65%)
Person of Color 98 (21.35%)
Age 30-49 233 (50.76%)
Age 18-29 104 (22.66%)
Age 50-64 79 (17.21%)
Age 65 or older 43 (9.37%)
Table 1. Descriptive statistics for our survey participants, N=459.
We also look at what types of compensation artists consider fair in exchange for the use of their artwork to train an AI art model. The results for this question can be seen in Table 3. We find that 50.97% of survey respondents stated that they did not need compensation, but what matters to them is who is profiting from the use of their artwork. The fair compensation option selected most by our survey participants was “I don’t need to profit, but I don’t want for-profit-companies to profit from my art” (22.80% selected this option), which indicates that artists are not wholesale against art being used in AI art models training data, but they oppose for-profit-companies reaping the financial benefits. Our results show that 36.56% of respondents did state they would find some type of compensation fair, only 1.29% of participants felt that a tax on companies and individuals who profit from the AI art outputs would be a fair option, and 11.18% of participants felt that they would not feel comfortable with any of the compensation options listed.
Question Strongly Disagree Disagree Neutral Agree Strongly Agree
AI threat to art workers 6.32% 14.16% 17.65% 33.77% 28.10%
AI positive development for art 13.07% 15.03% 27.02% 22.88% 22.00%
Required disclosure 2.83% 5.23% 11.98% 32.03% 48.15%
Output owned by model creators 32.46% 24.62% 18.74% 15.25% 11.55%
Output owned by AI user 19.61% 22.22% 21.57% 20.92% 18.30%
Output owned by the artist 9.37% 26.58% 25.05% 22.00% 19.39%
Table 2. Participant answers to our primary research questions by percent (answered on a 5-point Likert scale), N=459.
Descriptive statistics about our survey responses to demographic questions can be seen in Table 1. A full table of descriptive statistics that outlines participants who have used AI art models, their familiarity with Generative AI, professional and artistic status, whether or not they have purchased art, and art practices can be seen in Table 6.
Type Counts (%)
I don’t need to profit, but I don’t want for-profit-companies to profit from my art 106 (22.8%)
I would donate the use of my artwork to train AI art models (no compensation) 82 (17.63%)
A flat fee 55 (11.83%)
A portion of any profit made From the model creators 51 (10.97%)
Not comfortable with any listed options 42 (9.03%)
A portion of any profit from people who profit from derivative works made using the model 32 (6.88%)
A portion of any profit made 32 (6.88%)
I don’t need to profit, but I don’t want anyone else to profit from my art 27 (5.81%)
I don’t need to profit 22 (4.73%)
Other 10 (2.15%)
Tax 6 (1.29%)
Table 3. Counts percent per compensation type.
3.1. RQ1: Do artists see AI models as a threat to art workers?
Our results show that 61.87% of survey participants believe that AI models pose a threat to art workers. This finding is related to survey question no. 15 (seen in Section B). An ordinal logistic regression (OLR) was conducted to investigate the association between the Likert scale answers to RQ1 (dependent variable) and the secondary variables (as shown in Figure 3) in terms of the odds ratio of the OLR.
The results demonstrate a significant association (p-value ¡ 0.05, indicated in Table 4) between agreeing (represents a combination of both the responses ‘Agree’ and ‘Strongly Agree’ in the Likert) that AI models are a threat to art workers and age, as well as whether or not the participants have used AI to create art.
Figure 3. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ1 (AI models as a threat to art workers) and other survey variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ1 and the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 3. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ1 (AI models as a threat to art workers) and other survey variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Secondary Variable Participant Characteristic Primary Variable Likert Question Odds of Agreeing with RQ p-value
Used AI to make art RQ1: AI models are a threat to art workers 1.763 0.005
18-29 years old RQ1: AI models are a threat to art workers 2.329 0.002
Not makers RQ2: AI art models are a positive development in art 13.513 0.047
30-49 years old RQ2: AI art models are a positive development in art 1.535 0.039
Male and non-binary RQ2: AI art models are a positive development in art 1.663 0.007
Male RQ2: AI art models are a positive development in art 1.570 0.012
Not an artists RQ2: AI art models are a positive development in art 1.736 0.019
White RQ2: AI art models are a positive development in art 1.494 0.004
Used AI to make art RQ2: AI art models are a positive development in art 5.178 ¡0.001
Compensation: Would donate their - artwork to train models RQ2: AI art models are a positive development in art 15.444 ¡0.001
18-49 years old and ages above 65 required disclosure from model creators 2.439 0.001
18-29 years old RQ3: required disclosure from model creators 2.874 0.003
30-49 years old RQ3: required disclosure from model creators 1.777 0.028
Familiarity with AI art generally RQ3: required disclosure from model creators 2.620 0.003
Not an illustrator RQ3: required disclosure from model creators 4.822 0.026
Compensation: a portion of any - profit from derivatives RQ3: required to disclose 6.461 0.038
Compensation: portion - of any profit made RQ3: required disclosure from model creators 4.213 0.042
Has not purchased art RQ4_A: AI users owners of AI art made 1.773 0.039
Male RQ4_AA: users owners of AI art made 2.019 ¡0.001
Male and non-binary RQ4_A: I users owners of AI art made 2.197 ¡0.001
Compensation: did not select “not comfortable - with any listed compensation type” RQ4_A: users owners of AI art made 4.764 0.018
Person of color RQ4_B: Artists owners of AI art made 2.147 ¡0.001
Not used AI to make art RQ4_B:Artists owners of AI art made 1.523 0.0255
Person of color RQ4_C: Model creators owners of AI art made 2.491 ¡0.001
Male RQ4_C:Model creators owners of AI art made 2.033 ¡0.001
Male or non-binary RQ4_C: Model creators owners of AI art made 1.949 ¡0.001
Female or male RQ4_C: Model creators owners of AI art made 2.815 0.033
Not have a familiar with AI art generally RQ4_C: Model creators owners of AI art made 3.671 ¡0.001
Tattooists RQ4_C: Model creators owners of AI art made 6.674 0.042
Compensation: None, donate use - of artwork to train models RQ4_C: Model creators owners of AI art made 6.074 0.010
Table 4. Table represents the results of an ordinal logistic regression (OLR) to explore the relationship between our dependent variable (the Likert scale answers to each research question RQ1-4) and our secondary variables. The odds ratio is represented as the likelihood that the participants who selected a particular secondary variable agree or strongly agree with the primary variables research question. For example, for participants who have not used AI to make art, the odds of being more likely to agree that AI models are a threat to art workers is 1.7633 times that of participants who have used AI to make art (p-value=0.005). We consider evidence of a strong association to be a p-value less than 0.05.
3.2. RQ2: Do artists see AI art models as a positive development in the field of art?
We ran the same analysis for our second research question and found that 44.88% of the survey participants agree that AI art models are a positive development in the field of art. Notably, 27.02% of artists selected ’neutral’ on the Likert scale for this question. This analysis pertains to survey question number 14 (see the section on survey questions B). We ran an ordinal logistic regression (OLR) to explore the relationship between the Likert scale answers to RQ2 (the dependent variable) and the secondary variables (as seen in Figure 4, which shows the odds ratio of the OLR).
Our analysis indicates a strong association (p-value ¡ 0.05, as seen in Table 4) between agreeing that AI art models are a positive development in the field of art and various factors such as art practice, age, gender, artistic status, race, whether or not they have used AI to make art and compensation.
Figure 4. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ2 (AI models as a positive development) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the secondary variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ2 and the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 4. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ2 (AI models as a positive development) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
3.3. RQ3: What types of disclosure for the use of artwork to train an AI art model do artists consider fair?
We see from our survey results that 80.17% of the survey participants agree that creators of AI models should disclose the art and images used to train their models in detail. The analysis for RQ3 is related to survey question 13 (seen in Section B), which we used to run an ordinal logistic regression (OLR) to explore the relationship between the Likert scale answers to RQ3 and secondary variables. The results of the OLR, displayed in Figure 5 as the odds ratio, showed a strong association (p-value ¡ 0.05, as seen in Table 4) between agreeing with the disclosure requirement and factors such as art practice, age, familiarity with AI, whether or not they have used AI to create art and compensation preference.
Figure 5. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ3 (Should model creators be required to disclose in detail what art & images they use to train their AI models) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the secondary variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ3 and the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 5. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ3 (Should model creators be required to disclose in detail what art & images they use to train their AI models) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
3.4. RQ4: From an artist’s perspective, who should own the derivative artwork created by AI art models?
Finally, we look at the question that asks participants to think about a scenario where an AI art model was used by someone else to produce artwork recognizably in their (the survey participant’s) style and think about who should own the work and its derivatives. The analysis for RQ4 is related to survey questions B No. 8,9 and 10.
RQ4_A: If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh), Should that work and its derivatives be considered the property of the person who used the AI model to generate the artwork? We see that 39.22% of participants agreed that the work and its derivatives be considered the property of the person who used the AI model to generate the artwork. We run an ordinal logistic regression (OLR) to explore the relationship between our dependent variable (the Likert scale answers to RQ4_A) and our secondary variables, as seen in Figure 6 as the odds ratio of the OLR.
Figure 6. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_A (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) Should that work and its derivatives be considered the property of the person who used the AI model to generate the artwork?) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the secondary variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ4_Aand the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 6. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_A (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) Should that work and its derivatives be considered the property of the person who used the AI model to generate the artwork?) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
We see evidence of a strong association (p-value ¡ 0.05, seen in Table 4) between agreeing that the person who used the AI model to generate the artwork should be the owner, and gender, having purchased art, race, compensation, and having used AI to make art.
RQ4_B: If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh), Should that work and its derivatives be considered yours?
Our survey shows that 41.39% of the participants agreed that the work and its derivatives should be considered their property (the artist whose style is being recognizable and mimicked by the Generative AI art model). To explore the relationship between the Likert scale answers to RQ4_B and the secondary variables, we ran an ordinal logistic regression (OLR), as shown in Figure 7 with the odds ratio of the OLR. Table 4 shows that there is strong evidence of an association (p-value ¡ 0.05) between agreeing that the artist whose style is recognizable in the output should be the owner of the art and the participant’s race and whether they have used AI art models to create art.
Figure 7. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_B (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh) Should that work and its derivatives be considered yours? and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the secondary variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ4_B and the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 7. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_B (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh) Should that work and its derivatives be considered yours? and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
RQ4_C: If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh), Should that work and its derivatives be considered the property of the AI model creators?
Figure 8. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_C (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) Should that work and its derivatives be considered the property of the AI model creators?) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the secondary variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ4_C and the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 8. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_C (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) Should that work and its derivatives be considered the property of the AI model creators?) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
According to our survey results, only 26.80% of the participants believed that the work and its derivatives should be considered the property of the creators of AI models. To examine the relationship between the answers to RQ4_C on the Likert scale and other secondary variables, we conducted an ordinal logistic regression (OLR), as illustrated in Figure 8.
Our analysis suggests a strong association (with a p-value below 0.05, presented in Table 4) between agreement on AI companies owning the resulting artwork and art practice, race, gender, familiarity with AI art models, and compensation.
4. Discussion
The impact of Generative AI on artists is a complex global issue with impacts in many arenas, and it must be investigated through further multidisciplinary collaboration. We hope this work contributes to the existing literature on the ethics of AI-generated art and helps to provide a voice for artists on this topic.
In this study, we surveyed how artists feel about Generative Artificial Intelligence (AI) in art. Most participants believe that AI model creators should be required to disclose the art and images used to train them – an aspect of Generative AI tools that is currently far from transparent (see Sec. 2.4). However, only a minority of participants see AI art models as a positive development in art, while the majority see them as a threat to art workers.
When asked who should own the work and its derivatives when an AI art model is used to produce artwork in a particular artist’s style, the minority of participants believed that the AI model creators should own the work. In contrast, the majority of survey respondents were split between thinking it should belong to the person who used the AI model or the artist whose style is represented in the output.
Regarding compensation for using their artwork to train AI art models, over half of the participants did not require any compensation but were concerned about who was profiting from their artwork. Tentatively, we think this answer could reflect a belief within the artistic community: that art is the joint responsibility of all people (Huyghe, 1970). Corollary to that, there would be some onus on artists to share their work with each other and with the public, and there may be resistance to perceived exploitation of that practice, especially in the service of profit. We think this interpretation is consistent with the result that some participants stated that they do not require profit but do not want for-profit companies to profit from their art. Others considered some compensation fair, while a minority felt taxing companies and individuals who profit from AI art outputs would be a fair option. Finally, some participants stated they would feel uncomfortable with the compensation options listed. Understanding how the field of Generative AI falls short in respecting the values of the people who make the cultural artifacts it relies upon – and how that could be changed – is an important ethical question that we hope to expand on in future work.
In our future work, we plan to increase the sample size and expand it to a more representative sample. This will increase the statistical power of our study and also allow us to look at how opinions change over time. Moreover, we would like to put the current flourishing of Generative AI within the context of art history and taxonomy. We believe that a longer timeline would help consolidate, unify, and clarify the relationship between technology and art, which we only briefly touched upon in Section 2.5. While it is essential to understand what artists think about Generative AI, we need to collaborate with experts from various fields, including art, history, philosophy, psychology, cognitive science, and others, to understand the implications of Generative AI on our understanding of art and its role in our lives.
Acknowledgements.
Institutional Review Board Approval: The survey in this project is CHRBSS (Behavioral) STUDY00002379, approved by the University of Vermont I.R.B. on 3/14/2023. The authors would like to thank Maria Sckolnick, John Meluso, Kendall Fortney, Randall Harp, Laurent Hébert-Dufresne, Jean-Gabiel Young, and Winn Gillen, for their comments on early versions of this work. The Alfred P. Sloan Foundation, The National Science Foundation award #2242829, and MassMutual under the MassMutual Center of Excellence in Complex Systems and Data Science support this work. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the aforementioned financial supporters.
Appendix A Research Methods
A.1. Survey Methodology
Our anonymous original survey of artists is an observational study (N=516) of artists over the age of 18 years. To compare, we present a table of the distributions of the US Census art workforce 5. Incomplete surveys were dropped from the analysis, leaving a final N of 459.
Survey participants start by consenting to the IRB-approved survey information sheet and consent agreement. We then asked participants to answer several questions about their artistic identity and financial relationship with art. Before continuing, participants are asked to read a brief describing AI art and present an example. We then ask participants to disclose their familiarity with AI art models and tools. Then, we asked participants several questions about who they think “owns” art generated by AI models rated on a 5-point Likert scale. We ask several questions that invite artists to identify their ideal compensation criteria for their art being included in AI art models. We then ask participants to assess if AI art models are positive for their field, if they are a threat to their field, and if they think model creators should be required to disclose their data sources on a 5-point Likert scale. Lastly, we collect demographic information on the survey participants’ backgrounds.
Phase 1 Sampling: Survey participants for phase 1 (n = 252) of our observational study were recruited broadly on social media, using flyers locally at the University of Vermont, emailing artist groups, and direct contacts who are professional artists. Survey participants who are over the age of 18 and identify as an artist or work in an arts-related field were eligible. Phase 1 of the surveys was collected through a convenience sample from March 25, 2023, through September 8, 2023. Survey responses were collected on Qualtrics, an online survey platform.
Phase 2 Sampling: Phase 2 (n = 264) was collected via a targeted survey panel conducted by Qualtrics from September 12-23, 2023 on the Qualtrics online survey platform. A link to the complete survey and survey questions is available in Section B.
Descriptive statistics about the U.S. artist workforce are categorized as artists-related workers (US Census code 27-1010), which can be seen in Table 5. Data comes from the 2020 US Census, US Census Bureau, and ACS PUMS 5-Year Estimate. The occupational categories include Art Directors, Craft Artists, Fine Artists, Painters, Sculptors, Illustrators, Special Effects Artists and Animators, Artists and Related Workers, All Other Designers, Commercial and Industrial Designers, Fashion Designers, Floral Designers, Graphic Designers, Interior Designers, Merchandise Displayers and Window Trimmers, Set and Exhibit Designers, Designers, All Other. This data demonstrates the number of US workers potentially implicated by Generative AI art.
Type Number
U.S. artist workforce 144,000
Avg. Age 41
Avg. Salary \$57,958
Avg Male Salary 67,353
Avg Female Salary 45,618
Male 56.70%
Female 43.30%
White (Non-Hispanic) 73.20%
Asian (Non-Hispanic) 7.92%
Black 3.65%
Two or more 3.44%
White (Hispanic) 2.88%
American Indian 0.01%
Native Hawaiian 0.13%
Native other 0.04%
Table 5. Descriptive statistics about the US artist workforce categorized as artists-related workers (US Census code 27-1010). Data comes from the 2020 US Census, US Census Bureau, and ACS PUMS 5-Year Estimate.
A.2. Data
Original Data
Our original survey of artists is a convenience survey (N=459). For comparison of the distributions of the US art workforce) based on the demographics is presented in Table 5.
We transformed all survey response variables of interest into numerical form to analyze our survey results (survey questions 8,9,10,13,14,15 as seen in Section B). All Likert survey questions were converted from ‘Strongly Agree,’ ‘Agree,’ ‘Neutral,’ ‘Disagree,’ and ‘Strongly Disagree’ to an ordinal scale of 1,2,3,4,5. For our ordinal logistic regression analysis, we combine ‘Strongly Agree’ and ‘Agree’ to be equal to 3, ’Neutral’ to be equal to 2, and ‘Disagree’ and ‘Strongly Disagree’ to be equal to 1. In the original logistic regression, we reduced the dimensionality of the Likert scale to investigate the positive, neutral, and negative sentiment of the response questions in relation to our secondary variables.
All nominal and categorical variables (survey questions 1,2,3,4,5,11,12,16,17,18,19 as seen in Section B) were transformed into binary variables. Survey questions that included an “other” option or write-in answers were also categorized (e.g., in question 2, we added an additional art practice, which included digital art from the write-in answers).
Demographic questions (survey questions 16, 17, 18, 19 as seen in Section B) included asking survey participants their gender identity to identify whether they were white or a person of color, choosing their age range and country of residence. All demographic questions were transformed into binary variables.
A.3. Analytical Methods
Kruskal–Wallis Test The Kruskal-Wallis H-test is a non-parametric test that extends the Mann-Whitney U-test to compare differences between groups for an ordinal or continuous dependent variable. The explanatory groups may be ranked or unranked (Statistics, 2015a). Its null hypothesis is stochastic homogeneity, which is equivalent to equality of the expected value of rank means, conditional on the homogeneity of rank variances (Vargha and Delaney, 1998).
Cochran–Armitage Test The Cochran-Armitage test for trend uses weighted linear regression coefficients to test for trends in proportion between ordered levels of the response variable for a binary response variable. Its null hypothesis is a lack of trend, meaning that the proportions of each outcome are the same for each explanatory level (Iskander et al., 1983).
Odds ratio In general, the odds of an event are the probability of success over the probability of failure. This can be written
(1) $\frac{p}{1-p}$
where p is the probability of success. Unlike probability, odds can be from 1 to positive infinity  (Bruin, 2011a). The odds ratio (OR, or $\theta$) is simply the odds of one event divided by the odds of another. The OR can be from 0 to positive infinity, with an OR equal to 1 indicating equal odds of success between events. The natural logarithm of the OR is often used, as the transformation renders the distribution symmetric about 0  (Agresti, 2012).
Ordinal Logistic Regression Ordinal logistic regression (OLR) is a generalization of regression (either multiple linear or binomial logistic) used for an ordered categorical response variable and any number of categorical or numeric secondary variables (Statistics, 2015b). OLR finds the logit (also known as log-odds) of the dependent variable (Y) being less than or equal to a certain level (denoted j) (Bruin, 2011b). which can be written
(2) $logit(P(Y\leq j))=log\frac{P(Y\leq j)}{P(Y>j)}$
It is important to note that OLR assumes proportional odds, meaning the effect of the secondary variable(s) is the same between all levels of the dependent variable (Unit, 2016).
Appendix B Survey Questions
  1. (1)
    Would you consider yourself to be an artist? (options: Yes; Somewhat; No; I’m not sure)
  2. (2)
    What kind(s) of art do you practice? (select all that apply) (options: Photography; Paint; Drawing; Tattoo; Graphic Design; Illustration; Textile; Cinematography; Sculpture; Craft; Maker; Designer; Architecture; Jewelry; Writing; Music; Performance; Prefer not to respond; Not an artist but I work in an art-related field; Not an artist, but I am an art enthusiast; Not an artist; Other, please specify)
  3. (3)
    Have you ever purchased a piece of art? (options: Yes; No; I’m not sure)
  4. (4)
    Do you consider yourself to be a professional artist? (options: Yes; No; Prefer not to respond)
  5. (5)
    Please read the following description of Artificial Intelligence (AI) art before continuing: AI art is any computer-generated artwork created through the use of artificial intelligence. There are several ways to generate AI art. Models used to generate AI art: Artificial intelligence algorithms use two or more images to train (training is a process used where an algorithm is fed data to learn from) a model to generate new AI art (e.g., Lensa, DeepDream). Large language models use artificial intelligence, which uses large amounts of text data to train a model to generate new AI language (e.g., GPT-3). Text-to-image models use textual prompts (e.g., “a yellow elephant”) from a user to generate AI art that depicts the textual prompt. The AI art is trained on large amounts of image and text data to train the model and generate new AI art (e.g., DALL-E, Imagen). Example: Generated by DALL-E mini/Craiyon using the prompt “a cat in the style of Van Gogh.” According to its creators, “[t]he model was trained on unfiltered data from the Internet, limited to pictures with English descriptions.” Includes image seen in Figure 1.
  6. (6)
    Which method of AI art listed below had you heard of before starting this survey? (select all that apply) (options: AI art generally; Artificial intelligence or deep learning algorithms, such as generative adversarial networks (GANs); Large Language Models; Text-to-image models: None of the above)
  7. (7)
    Have you ever used AI to generate art? (options: Yes; No; I don’t know)
  8. (8)
    If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh), Should that work and its derivatives be considered the property of the person who used the AI model to generate the artwork? (options: Strongly agree; Agree; Neutral; Disagree; Strongly disagree)
  9. (9)
    If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh), Should that work and its derivatives be considered yours? (options: Strongly agree; Agree; Neutral; Disagree; Strongly disagree)
  10. (10)
    If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh), Should that work and its derivatives be considered the property of the AI model creators? (options: Strongly agree; Agree; Neutral; Disagree; Strongly disagree)
  11. (11)
    What type of compensation for the use of your artwork to train an AI art model would you consider fair? (select all that apply) (options: I would donate the use of my artwork to train AI art models (no compensation); I don’t need to profit; I don’t need to profit, but I don’t want for-profit-companies to profit from my art; I don’t need to profit, but I don’t want anyone else to profit from my art; A portion of any profit made; A portion of any profit made From the model creators; A portion of any profit from people who profit from derivative works made using the model; A flat fee; A flat fee from the model creators, for the use of my artwork as training data; flat fee for each derivative work sold (by anyone using the model); companies and individuals should be subject to a new tax by the government if they profit from AI art trained on other people’s work; I would not feel comfortable with any of the options listed above; Other, please specify)
  12. (12)
    You selected the response: Companies and individuals should be subject to a new tax in order to use AI art that was trained with your artwork. In your opinion, what should these taxes support? (option: open text field)
  13. (13)
    Do you think model creators should be required to disclose in detail what art and images they use to train their AI models? (options: Strongly agree; Agree; Neutral; Disagree; Strongly disagree)
  14. (14)
    Do you agree with the following statement? ”I see AI art models as a positive development in the field of art.” (options: Strongly agree; Agree; Neutral; Disagree; Strongly disagree)
  15. (15)
    Do you agree with the following statement? ”I see AI models as a threat to art workers?” (options: Strongly agree; Agree; Neutral; Disagree; Strongly disagree)
  16. (16)
    What is your age (options: 18-29; 30-49; 50-64, 65 or older)
  17. (17)
    Do you identify as a person of color? (options: Yes, No)
  18. (18)
    What is your gender identity? (options: Non-binary; Female; Male)
  19. (19)
    In which country do you currently reside? (options: drop-down menu of all countries)
Appendix C Full table of descriptive statistics
A complete table of descriptive statistics can be found in Table 6.
Type Count %
Male 273 (59.48%)
Female 163 (35.51%)
Non-binary 23 (5.01%)
White 361 (78.65%)
Person of Color 98 (21.35%)
Age 30-49 233 (50.76%)
Age 18-29 104 (22.66%)
Age 50-64 79 (17.21%)
Age 65 or older 43 (9.37%)
Used AI Models - No 237 (49.17%)
Used AI Models - Yes 231 (47.93%)
Used AI Models - I don’t know 10 (2.07%)
Used AI Models - None of the above 4 (0.83%)
Familiarity - AI art generally 379 (78.63%)
Familiarity - AI/Deeplearning/GANS 42 (8.71%)
Familiarity - Text-to-image 12 (2.49%)
Familiarity - LLMS 5 (1.04%)
Familiarity - None 44 (9.13%)
Professional Artist - No 256 (52.57%)
Professional Artist - Yes 214 (43.94%)
Professional Artist - Prefer not to respond 17 (3.49%)
Artist - Yes 310 (62.63%)
Artist - Somewhat 176 (35.56%)
Artist - No 7 (1.41%)
Artist - I’m not sure 2 (0.40%)
Purchased Art - Yes 422 (86.65%)
Purchased Art - No 59 (12.11%)
Purchased Art - I’m not sure 6 (1.23%)
Art Practice - Photography 116 (23.58%)
Art Practice - Paint 167 (33.94%)
Art Practice - Writing 52 (10.57%)
Art Practice - Designer 43 (8.74%)
Art Practice - Craft 20 (4.07%)
Art Practice - Music 18 (3.66%)
Art Practice - Digital 14(2.85%)
Art Practice - Not artist 13(2.64%)
Art Practice - Other 11(2.24%)
Art Practice - Illustration 9 (1.83%)
Art Practice - Tattoo 9 (1.83%)
Art Practice - Drawing 8 (1.63%)
Art Practice - Sculpture 7 (1.42%)
Art Practice - Maker 5 (1.02%)
Table 6. Descriptive statistics for our survey participants.
Appendix D All tables for ordinal logistic regressions
We ran an ordinal logistic regression (OLR) to explore the relationship between our dependent variable (the Likert scale answers to each primary research question). Our secondary variables, the results of these original logistic regressions, their odds-ratios, their p-values, and their confidence intervals (CI) can be seen in Tables 7, 8, 9, 10, 11, 12.
Variable CI 2.5% 97.5% Odds Ratio p-value
Artist 0.364481 0.913316 0.576963 ¡0.001
Purchased art 0.331723 1.004549 0.577263 ¡0.001
Person of color 1.23678 3.085823 1.953582 ¡0.001
Professional artist 0.840948 2.063429 1.317284 ¡0.001
Used AI art models 3.4805 7.703421 5.178007 ¡0.001
Male 1.105324 2.232347 1.570817 0.011789
Female 0.417108 0.868241 0.601789 ¡0.001
Non-binary 0.420314 2.076117 0.934141 ¡0.001
Age 50-64 0.5506 1.369427 0.868336 ¡0.001
Age 65 plus 0.346227 1.0817 0.611975 ¡0.001
Age 18-29 0.56397 1.493795 0.917854 0.730133
Age 30-49 1.022423 2.304733 1.535061 0.038743
Familiarity AI art gen 0.594964 1.937282 1.073598 0.813584
Familiarity AI/DL 0.460722 2.345589 1.039551 0.925566
Familiarity LLM 0.185243 8.879363 1.282512 0.801008
Familiarity text to image 0.256932 3.219854 0.90955 0.883141
Photographer 0.267789 3.76309 1.003849 0.995454
Crafter 0.059216 1.147651 0.260691 0.075429
Designer 0.203683 3.386843 0.830568 0.795733
Digital Artist 0.126946 2.668446 0.582021 0.486019
Sketch artist 0.222575 9.567887 1.459305 0.693624
Illustrator 0.100764 3.509196 0.594644 0.566036
Musician 0.11112 2.489955 0.526007 0.417996
Maker 0.005815 0.967095 0.074991 0.047077
Sculptor 0.016156 1.091709 0.132805 0.060336
Other artist 0.100764 3.509204 0.594644 0.566036
Painter 0.188397 2.563246 0.694917 0.5847
Tattoist 0.269396 11.050416 1.725382 0.564828
Writer 0.095046 1.50597 0.378334 0.167879
None, don’t need profit 0.779115 12.922381 3.173015 0.107049
None, donate use of artwork to train models 4.162129 57.308125 15.444217 ¡0.001
Flat fee 0.445466 5.012977 1.49436 0.515373
None, but don’t want for-profit comp to profit 0.305542 3.134003 0.978555 0.970882
None, but don’t want anyone else to profit 0.387164 5.704167 1.486085 0.563771
Not comfortable with any listed 0.0935 1.160233 0.329365 0.083874
Portion of any profit from derivatives 0.296462 3.932296 1.079711 0.907418
Portion of any profit from model creators 0.295023 3.38909 0.999929 0.999909
Portion of any profit 0.319872 4.208076 1.160191 0.821179
Tax 0.327069 13.554205 2.105507 0.433234
Table 7. Ordinal logistic regression (OLR). Dependent variable: agreeing or strongly agreeing that AI is a positive development for the field of art. Strong evidence follows a threshold of a p-value less than 0.05.
Variable CI 2.5% CI 97.5% Odds Ratio p-value
Artist 0.673062 1.734431 1.080453 ¡0.001
Purchased art 0.56515 1.850283 1.022589 ¡0.001
Person of color 0.835831 2.157276 1.342802 ¡0.001
Professional artist 0.511389 1.27991 0.809031 ¡0.001
Used AI art models 0.381466 0.843053 0.567094 ¡0.001
Male 0.68315 1.452222 0.996035 ¡0.001
Female 0.678301 1.484131 1.003338 ¡0.001
Non-binary 0.41613 2.447633 1.009224 ¡0.001
Age 50-64 0.399295 1.029979 0.641299 ¡0.001
Age 65 plus 0.385716 1.286586 0.704455 ¡0.001
Age 18-29 1.337791 4.055724 2.329316 ¡0.001
Age 30-49 0.829556 1.938754 1.26819 ¡0.001
Familiarity AI art gen 0.55586 2.015918 1.058569 0.862508
Familiarity AI/DL 0.644924 4.062339 1.61861 0.305025
Familiarity LLM 0.160508 6.218764 0.999079 0.999212
Familiarity text to image 0.143725 1.825538 0.512226 0.302196
Photographer 0.624162 8.7873 2.341944 0.207189
Crafter 0.691218 18.224995 3.549288 0.129129
Designer 0.625025 10.555753 2.568582 0.190796
Digital Artist 0.349361 7.951837 1.666752 0.521639
Sketch artist 0.274778 13.110306 1.898007 0.51577
Illustrator 0.35758 15.847783 2.380514 0.369873
Musician 0.845339 24.983415 4.59559 0.077486
Maker 0.419635 68.154064 5.347879 0.196624
Sculptor 0.274778 13.110325 1.898006 0.515771
Other artist 0.603302 34.971648 4.593306 0.141009
Painter 0.716997 9.797998 2.650498 0.14395
Tattoist 0.603301 34.971718 4.593306 0.141009
Writer 0.58997 9.564791 2.375487 0.223434
None, don’t need profit 0.318227 4.551266 1.203467 0.784937
None, donate use of artwork to train models 0.366867 3.7479 1.172597 0.788264
Flat fee 0.492379 5.558458 1.654347 0.415569
None, but don’t want for-profit comp to profit 0.672149 6.774345 2.133862 0.198466
None, but don’t want anyone else to profit 0.427385 6.406989 1.654767 0.465872
Not comfortable with any listed 0.880466 11.318314 3.156802 0.077639
Portion of any profit from derivatives 0.76601 11.308242 2.943165 0.11599
Portion of any profit from model creators 0.690259 8.054443 2.357891 0.171143
Portion of any profit 0.416284 5.403195 1.499755 0.535396
Tax 0.131044 3.68014 0.69445 0.668239
Table 8. Ordinal logistic regression (OLR). Dependent variable: agreeing or strongly agreeing that AI models are a threat to art workers. Strong evidence follows a threshold of a p-value less than 0.05.
Variable CI 2.5% CI 97.5% Odds Ratio p-value
Artist 0.618355 1.949844 1.098041 ¡0.001
Purchased art 0.851732 3.164821 1.641822 ¡0.001
Person of color 0.809301 2.740342 1.489215 ¡0.001
Professional artist 0.640443 2.00734 1.133837 ¡0.001
Used AI art models 0.38645 1.045068 0.635505 ¡0.001
Male 0.587018 1.506886 0.940515 ¡0.001
Female 0.669859 1.797992 1.097452 ¡0.001
Non-binary 0.30543 2.406881 0.857399 ¡0.001
Age 50-64 0.236998 0.712557 0.410944 ¡0.001
Age 65 plus 0.325902 1.57366 0.716142 ¡0.001
Age 18-29 1.427261 5.790101 2.874715 ¡0.001
Age 30-49 1.061875 2.97578 1.777612 ¡0.001
Familiarity AI art gen 1.363085 5.05368 2.624614 ¡0.001
Familiarity AI/DL 0.879414 6.605245 2.410134 ¡0.001
Familiarity text to image 0.131881 1.59315 0.458373 ¡0.001
Photographer 0.026665 1.656513 0.21017 0.138655
Crafter 0.020146 1.971367 0.199289 0.167743
Designer 0.021091 1.532809 0.179801 0.116568
Digital Artist 0.013077 1.343422 0.132543 0.087241
Sketch artist 0.018934 6.359606 0.347009 0.475682
Illustrator 0.006274 0.725497 0.067464 0.026098
Musician 0.022443 2.549968 0.239228 0.23615
Other artist 0.01451 2.365917 0.185285 0.194532
Painter 0.036576 2.225627 0.285316 0.231452
Tattoist 0.022214 7.300981 0.402718 0.538408
Writer 0.017016 1.172106 0.141224 0.069849
None, don’t need profit 0.105173 1.407149 0.384701 0.14881
None, donate use of artwork to train models 0.446605 4.575144 1.429434 0.547225
Flat fee 0.670956 8.716645 2.418365 0.177032
None, but don’t want for-profit comp to profit 0.639246 6.378818 2.019314 0.231116
None, but don’t want anyone else to profit 0.643891 15.548169 3.164068 0.156186
Not comfortable with any listed 0.559388 7.355529 2.028445 0.281879
Portion of any profit from derivatives 1.100951 37.925791 6.461768 0.038783
Portion of any profit from model creators 1.051359 16.889636 4.213914 0.042288
Portion of any profit 0.327584 4.283127 1.184518 0.796243
Table 9. Ordinal logistic regression (OLR). Dependent variable: agreeing or strongly agreeing that model creators should be required to disclose in detail what art & images they use to train their AI models. Strong evidence follows a threshold of a p-value less than 0.05.
Variable CI 2.5% CI 97.5% Odds Ratio p-value
Artist 0.468541 1.160369 0.737347 0.187823
Purchased art 0.439453 1.302365 0.756523 0.314052
Person of color 1.603519 3.87071 2.491336 ¡0.001
Professional artist 0.539563 1.334183 0.848455 0.476731
Used AI art models 0.684392 1.486553 1.008655 0.965265
Male 1.394213 2.965517 2.033362 ¡0.001
Female 0.34673 0.759067 0.513022 ¡0.001
Non-binary 0.136972 0.921049 0.355187 0.033244
Age 50-64 0.547054 1.443719 0.888703 0.633636
Age 65 plus 0.621558 2.010455 1.117861 0.709854
Age 18-29 0.488074 1.356759 0.813756 0.429419
Age 30-49 0.750917 1.753871 1.147611 0.524626
Familiarity AI art generally 0.152611 0.486134 0.272377 ¡0.001
Familiarity AI/DL 0.375901 1.823414 0.827903 0.639206
Familiarity LLM 0.120181 3.748762 0.671215 0.649641
Familiarity text to image 0.140257 1.564695 0.468465 0.217808
Photographer 0.487757 6.18452 1.73682 0.394225
Crafter 0.174597 3.568943 0.789385 0.758671
Designer 0.207425 3.227255 0.818176 0.774409
Digital Artist 0.037919 1.787102 0.260319 0.170916
Sketch artist 0.289616 10.860348 1.773507 0.535461
Illustrator 0.239963 8.326074 1.41349 0.702103
Musician 0.256817 5.623486 1.201752 0.815437
Maker 0.035642 5.508235 0.443085 0.526711
Sculptor 0.121091 8.227352 0.998126 0.998609
Other artist 0.239959 8.326229 1.41349 0.702106
Painter 0.4663 5.715676 1.632549 0.443291
Tattoist 1.065367 41.810186 6.674068 0.042601
Writer 0.088428 1.500551 0.364267 0.16208
None, don’t need profit 0.76008 16.216441 3.510811 0.107708
None, donate use of artwork to train models 1.512847 24.393075 6.074783 0.01097
Flat fee 0.871625 14.630981 3.571096 0.076892
None, but don’t want for-profit comp to profit 0.30342 4.788926 1.205429 0.790657
None, but don’t want anyone else to profit 0.177118 4.423683 0.885163 0.881873
Not comfortable with any listed 0.215123 4.09741 0.938854 0.933113
Portion of any profit from derivatives 0.285511 5.930689 1.30126 0.733653
Portion of any profit from model creators 0.292329 5.254805 1.239407 0.770881
Portion of any profit 0.621704 11.541633 2.678708 0.186106
Tax 0.355144 20.211859 2.679203 0.339136
Table 10. Ordinal logistic regression (OLR). Dependent variable: agreeing or strongly agreeing that if an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) that work and its derivatives should be considered the property of the AI model creators. Strong evidence follows a threshold of a p-value less than 0.05.
Variable CI 2.5% CI 97.5% Odds Ratio p-value
Artist 0.601364 1.425504 0.925876 0.726499
Purchased art 0.682946 1.985811 1.164561 0.57583
Person of color 1.382372 3.336151 2.147511 ¡0.001
Professional artist 0.857238 2.014212 1.314024 0.210158
Used AI art models 0.453536 0.949834 0.656341 0.025556
Male 0.929958 1.863216 1.316326 0.121057
Female 0.540109 1.114689 0.775921 ¡0.001
Non-binary 0.304406 1.426784 0.659031 ¡0.001
Age 50-64 0.433393 1.072792 0.681865 ¡0.001
Age 65 plus 0.715666 2.230618 1.263478 ¡0.001
Age 18-29 0.749478 1.976648 1.21715 0.427012
Age 30-49 0.76602 1.711638 1.145054 0.508996
Familiarity AI art gen 0.341862 1.085694 0.609227 0.092752
Familiarity AI/DL 0.454393 2.193558 0.998367 0.996753
Familiarity LLM 0.239351 9.063794 1.472899 0.676175
Familiarity text to image 0.199754 2.904406 0.761687 0.690169
Photographer 0.767339 8.542394 2.560257 0.126216
Crafter 0.330235 5.508344 1.348721 0.6769
Designer 0.350222 4.596068 1.268718 0.71705
Digital Artist 0.350639 7.202454 1.58917 0.547998
Sketch artist 0.442624 12.710739 2.371937 0.313265
Illustrator 0.705023 21.346164 3.879373 0.119181
Musician 0.846183 16.829236 3.773675 0.081682
Maker 0.093845 6.014805 0.751304 0.787605
Sculptor 0.563212 23.470174 3.635751 0.174907
Other artist 0.423624 14.806303 2.504456 0.311247
Painter 0.742785 7.928239 2.426721 0.142187
Tattoist 0.705023 21.346176 3.879373 0.119182
Writer 0.37674 4.75844 1.338916 0.65191
None, don’t need profit 0.25232 3.961372 0.999766 0.999734
None, donate use of artwork to train models 0.513752 5.733981 1.716346 0.380072
Flat fee 0.358686 4.194022 1.226515 0.744814
None, but don’t want for-profit comp to profit 0.334713 3.531268 1.08718 0.889399
None, but don’t want anyone else to profit 0.418543 6.478014 1.646611 0.475448
Not comfortable with any listed 0.349148 4.307525 1.226361 0.750228
Portion of any profit from derivatives 0.409981 5.41326 1.489744 0.544844
Portion of any profit from model creators 0.573956 6.71113 1.962624 0.282421
Portion of any profit 0.435868 5.718306 1.578742 0.486822
Tax 0.532577 29.762374 3.981302 0.178264
Table 11. Ordinal logistic regression (OLR). Dependent variable: agreeing or strongly agreeing that if an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh) that work and its derivatives should be considered the property of the artist. Strong evidence follows a threshold of a p-value less than 0.05.
Variable CI 2.5% CI 97.5% Odds Ratio p-value
Artist 0.414702 1.019737 0.650297 ¡0.001
Purchased art 0.327522 0.974205 0.564866 ¡0.001
Person of color 1.120118 2.725496 1.747248 ¡0.001
Professional artist 0.702605 1.699884 1.092862 ¡0.001
Used AI art models 2.034051 4.377637 2.984013 ¡0.001
Male 1.41509 2.883435 2.01998 ¡0.001
Female 0.314178 0.66174 0.455965 ¡0.001
Non-binary 0.389118 1.943853 0.869706 ¡0.001
Age 50-64 0.698376 1.745508 1.104093 0.671752
Age 65 plus 0.352683 1.146873 0.63599 0.132472
Age 18-29 0.570787 1.525004 0.93298 0.782005
Age 30-49 0.791308 1.788877 1.18977 0.40368
Familiarity AI art generally 0.301199 1.004236 0.549977 0.05163
Familiarity AI/DL 0.388081 1.979027 0.876369 0.750841
Familiarity LLM 0.086738 2.268525 0.443584 0.328954
Familiarity text to image 0.221942 2.714057 0.776121 0.691518
Photographer 0.460161 4.790509 1.484723 0.50843
Crafter 0.116186 1.930897 0.473649 0.297291
Designer 0.204141 2.595388 0.727891 0.624392
Digital Artist 0.174549 3.434168 0.77423 0.736364
Sketch artist 0.132029 4.012481 0.727848 0.715319
Illustrator 0.480708 17.524118 2.902409 0.245438
Musician 0.237822 4.606483 1.046672 0.95189
Maker 0.065305 3.981952 0.509941 0.520716
Sculptor 0.15958 5.25934 0.916126 0.921737
Other artist 0.224562 6.596202 1.21707 0.819786
Painter 0.432178 4.332419 1.368348 0.593818
Tattoist 0.554963 15.175121 2.902004 0.206845
Writer 0.230251 2.832793 0.807621 0.738605
None, don’t need profit 0.498665 8.015598 1.999275 0.328153
None, donate use of artwork to train models 0.965801 11.105517 3.275014 0.056896
Flat fee 0.259222 3.002946 0.882286 0.841163
None, but don’t want for-profit comp to profit 0.159296 1.686612 0.518334 0.274998
None, but don’t want anyone else to profit 0.126867 1.973433 0.500362 0.322656
Not comfortable with any listed 0.05761 0.765249 0.209967 0.018008
Portion of any profit from derivatives 0.164151 2.258827 0.608925 0.458287
Portion of any profit from model creators 0.108502 1.319073 0.378315 0.127163
Portion of any profit 0.315746 4.308758 1.166393 0.817425
Tax 0.132833 5.021371 0.816703 0.827032
Table 12. Ordinal logistic regression (OLR). Dependent variable: agreeing or strongly agreeing that if an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) that work and its derivatives should be considered the property of the person who used the AI model to generate the artwork. Strong evidence follows a threshold of a p-value less than 0.05.
Appendix E Contingency Tables:
All contingency tables for dependent variables with secondary variables can be seen in Tables 13, 14, 15, 16, 18, 19, 20, 21, 17, 22, 23, 24, and 25.
Secondary Variable (Binary) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Not Person of color AI Positive for Art [[50 53 90 76 66]
Person of color AI Positive for Art [ 6 11 22 24 29]]
Not Person of color AI Threat to Art [[ 22 53 54 113 93]
Person of color AI Threat to Art [ 6 10 14 32 30]]
Not Person of color Required Disclosure [[ 9 21 38 97 170]
Person of color Required Disclosure [ 3 2 10 37 40]]
Not Person of color Owner AI Company [[124 79 59 41 32]
Person of color Owner AI Company [ 18 19 14 23 18]]
Not Person of color Owner AI User [[75 67 73 61 59]
Person of color Owner AI User [ 8 23 13 27 21]]
Not Person of color Owner Artist [[36 94 81 68 56]
Person of color Owner Artist [ 5 17 18 24 28]]
Table 13. Contingency table for the secondary variable of interest (person of color) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (Binary) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Not Professional Artist AI Positive for Art [[33 34 63 60 39]
Professional Artist AI Positive for Art [23 30 49 40 56]]
Not Professional Artist AI Threat to Art [[12 31 35 89 62]
Professional Artist AI Threat to Art [16 32 33 56 61]]
Not Professional Artist Required Disclosure [[ 5 11 31 89 93]
Professional Artist Required Disclosure [7 12 17 45 117]]
Not Professional Artist Owner AI Company [[59 56 47 45 22]
Professional Artist Owner AI Company [83 42 26 19 28]]
Not Professional Artist Owner AI User [[37 52 53 52 35]
Professional Artist Owner AI User [46 38 33 36 45]]
Not Professional Artist Owner Artist [[15 67 59 50 38]
Professional Artist Owner Artist [26 44 40 42 46]]
Table 14. Contingency table for the secondary variable of interest (professional artist) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (Binary) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Not Purchased Art AI Positive for Art [[ 7 3 17 12 10]
Purchased Art AI Positive for Art [49 61 95 88 85]]
Not Purchased Art AI Threat to Art [[ 1 6 9 21 12]
Purchased Art AI Threat to Art [27 57 59 124 111]]
Not Purchased Art Required Disclosure [[ 0 1 13 19 16]
Purchased Art Required Disclosure [12 22 35 115 194]]
Not Purchased Art Owner AI Company [[ 10 12 14 7 6]
Purchased Art Owner AI Company [132 86 59 57 44]]
Not Purchased Art Owner AI User [[ 4 11 12 13 9]
Purchased Art Owner AI User [79 79 74 75 71]]
Not Purchased Art Owner Artist [[ 4 13 13 12 7]
Purchased Art Owner Artist [37 98 86 80 77]]
Table 15. Contingency table for the secondary variable of interest (purchased art) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (Binary) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Not Used AI art AI Positive for Art [[44 45 65 38 19]
Used AI art AI Positive for Art [12 19 47 62 76]]
Not Used AI art AI Threat to Art [[ 6 24 35 76 70]
Used AI art AI Threat to Art [22 39 33 69 53]]
Not Used AI art Required Disclosure [[ 4 8 24 64 111]
Used AI art Required Disclosure [ 8 15 24 70 99]]
Not Used AI art Owner AI Company [[58 54 47 36 16]
Used AI art Owner AI Company [84 44 26 28 34]]
Not Used AI art Owner AI User [[49 58 45 29 30]
Used AI art Owner AI User [34 32 41 59 50]]
Not Used AI art Owner Artist [[17 46 56 48 44]
Used AI art Owner Artist [24 65 43 44 40]]
Table 16. Contingency table for the secondary variable of interest (used AI art) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (binary) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Not Artist Owner AI User [[ 1 0 4 1 0]
Artist Owner AI User [88 102 94 95 84]]
Not Artist Owner Artist [[ 0 5 0 0 1]
Artist Owner Artist [43 116 115 101 87]]
Not Artist Owner AI Company [[ 3 0 2 1 0]
Artist Owner AI Company [145 113 84 68 53]]
Not Artist Required Disclosure [[ 0 0 0 3 2]
Artist Required Disclosure [13 24 55 143 218]]
Not Artist AI Positive for Art [[ 1 0 1 3 0]
Artist AI Positive for Art [59 69 123 101 100]]
Not Artist AI Threat to Art [[ 0 1 2 1 1]
Artist AI Threat to Art [28 64 78 154 128]]
Table 17. Contingency table for the secondary variable of interest (Artist) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (Categorical) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Age 18-29 Owner AI User [[42 46 46 57 40]
Age 30-49 Owner AI User [24 24 17 22 17]
Age 50-64 Owner AI User [16 12 19 11 21]
Age 65 plus Owner AI User [ 5 16 12 4 6]]
Age 18-29 Owner Artist [[21 62 49 61 38]
Age 30-49 Owner Artist [10 27 20 18 29]
Age 50-64 Owner Artist [10 21 24 10 14]
Age 65 plus Owner Artist [ 2 6 17 11 6]]
Age 18-29 Owner AI Company [[77 47 36 38 33]
Age 30-49 Owner AI Company [33 30 19 12 10]
Age50-64 Owner AI Company [29 16 14 12 8]
Age 65 plus Owner AI Company [ 6 15 13 7 2]]
Age 18-29 Required Disclosure [[ 6 13 25 75 112]
Age 30-49 Required Disclosure [ 1 2 10 38 53]
Age 50-64 Required Disclosure [ 6 5 14 19 35]
Age 65 plus Required Disclosure [ 0 4 5 14 20]]
Age 18-29 AI Positive for Art [[22 33 60 59 56]
Age 30-49 AI Positive for Art [20 19 22 20 23]
Age 50-64 AI Positive for Art [12 9 25 15 18]
Age 65 plus AI Positive for Art [ 6 8 17 9 3]]
Age 18-29 AI Threat to Art [[16 37 35 78 65]
Age 30-49 AI Threat to Art [ 7 9 10 44 34]
Age 50-64 AI Threat to Art [ 4 13 21 22 18]
Age 65 plus AI Threat to Art [ 1 6 13 11 12]]
Table 18. Contingency table for the secondary variable of interest (Age) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (Categorical) Dependent Variable (Likert) Contingency Table [SD D N A SA]
AI Model Familiarity: general Owner AI User [[82 81 67 73 67]
AI Model Familiarity: deep learning Owner AI User [ 2 8 14 9 8]
AI Model Familiarity: large language models Owner AI User [ 5 8 11 11 7]
AI Model Familiarity: text-to-image Owner AI User [ 0 3 4 3 1]
AI Model Familiarity: none Owner AI User [ 0 2 2 0 1]]
AI Model Familiarity: general Owner Artist [[ 40 106 76 80 68]
AI Model Familiarity: deep learning Owner Artist [ 1 3 20 8 8]
AI Model Familiarity: large language models Owner Artist [ 2 7 16 8 9]
AI Model Familiarity: text-to-image Owner Artist [ 0 4 2 2 3]
AI Model Familiarity: none Owner Artist [ 0 1 1 3 0]]
AI Model Familiarity: general Owner AI Company [[137 97 49 46 41]
AI Model Familiarity: deep learning Owner AI Company [ 3 3 20 11 4]
AI Model Familiarity: large language models Owner AI Company [ 7 8 11 10 6]
AI Model Familiarity: text-to-image Owner AI Company [ 1 3 5 1 1]
AI Model Familiarity: none Owner AI Company [ 0 2 1 1 1]]
AI Model Familiarity: general Required Disclosure [[ 10 18 34 109 193]
AI Model Familiarity: deep learning Required Disclosure [ 1 3 12 14 10]
AI Model Familiarity: large language models Required Disclosure [ 1 2 5 18 13]
AI Model Familiarity: text-to-image Required Disclosure [ 1 1 4 2 2]
AI Model Familiarity: none Required Disclosure [ 0 0 0 3 2]]
AI Model Familiarity: general AI Positive for Art [[52 51 94 82 84]
AI Model Familiarity: deep learning AI Positive for Art [ 4 5 17 9 5]
AI Model Familiarity: large language models AI Positive for Art [ 4 8 10 10 7]
AI Model Familiarity: text-to-image AI Positive for Art [ 0 3 3 1 3]
AI Model Familiarity: none AI Positive for Art [ 0 2 0 2 1]]
AI Model Familiarity: general AI Threat to Art [[ 26 51 59 117 110]
AI Model Familiarity: deep learning AI Threat to Art [ 0 6 11 17 6]
AI Model Familiarity: large language models AI Threat to Art [ 1 5 6 16 11]
AI Model Familiarity: text-to-image AI Threat to Art [ 1 2 3 3 1]
AI Model Familiarity: none AI Threat to Art [ 0 1 1 2 1]]
Table 19. Contingency table for the secondary variable of interest (AI model familiarity) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (Categorical) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Art Practice: Paint Owner AI User [[28 30 36 41 28]
Art Practice: Photography Owner AI User [11 29 22 27 23]
Art Practice: Writing Owner AI User [14 12 6 6 11]
Art Practice: Designer Owner AI User [11 12 4 7 7]
Art Practice: Craft Owner AI User [ 4 6 7 1 0]
Art Practice: Music Owner AI User [ 5 4 1 3 4]
Art Practice: Digital Owner AI User [ 4 3 3 2 2]
Art Practice: Not Artist Owner AI User [ 2 2 5 1 1]
Art Practice: Other Owner AI User [ 4 0 2 4 0]
Art Practice: Tattoo Owner AI User [ 0 0 4 2 2]
Art Practice: Illustration Owner AI User [ 3 0 0 2 4]
Art Practice: Drawing Owner AI User [ 0 3 3 0 1]
Art Practice: Sculpture Owner AI User [ 1 1 3 0 1]
Art Practice Maker Owner AI User [ 2 0 2 0 0]]
Art Practice: Paint Owner Artist [[13 38 41 37 34]
Art Practice: Photography Owner Artist [ 6 30 22 30 23]
Art Practice: Writing Owner Artist [ 7 15 12 8 7]
Art Practice: Designer Owner Artist [ 9 12 8 7 5]
Art Practice: Craft Owner Artist [ 1 6 7 2 2]
Art Practice: Music Owner Artist [ 1 3 4 7 2]
Art Practice: Digital Owner Artist [ 1 5 3 4 1]
Art Practice: Not Artist Owner Artist [ 0 6 4 0 1]
Art Practice: Other Owner Artist [ 2 2 1 3 2]
Art Practice: Tattoo Owner Artist [ 0 1 3 0 4]
Art Practice: Illustration Owner Artist [ 1 1 2 1 4]
Art Practice: Drawing Owner Artist [ 0 1 4 1 1]
Art Practice: Sculpture Owner Artist [ 0 1 2 1 2]
Art Practice Maker Owner Artist [ 2 0 2 0 0]]
Art Practice: Paint Owner AI Company [[42 39 29 28 25]
Art Practice: Photography Owner AI Company [30 25 20 22 15]
Art Practice: Writing Owner AI Company [25 14 7 2 1]
Art Practice: Designer Owner AI Company [17 10 6 5 3]
Art Practice: Craft Owner AI Company [ 6 5 6 1 0]
Art Practice: Music Owner AI Company [ 3 7 3 2 2]
Art Practice: Digital Owner AI Company [ 6 6 1 1 0]
Art Practice: Not Artist Owner AI Company [ 4 2 4 1 0]
Art Practice: Other Owner AI Company [ 4 0 4 2 0]
Art Practice: Tattoo Owner AI Company [ 1 0 2 1 4]
Art Practice: Illustration Owner AI Company [ 3 2 1 1 2]
Art Practice: Drawing Owner AI Company [ 1 2 2 2 0]
Art Practice: Sculpture Owner AI Company [ 4 0 0 1 1]
Art Practice Maker Owner AI Company [ 2 1 1 0 0]]
Table 20. Contingency table for the secondary variable of interest (Art Practice) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (Categorical) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Art Practice: Paint Required Disclosure [[ 2 5 20 64 71]
Art Practice: Photography Required Disclosure [ 2 7 13 28 59]
Art Practice: Writing Required Disclosure [ 5 3 4 14 20]
Art Practice: Designer Required Disclosure [ 3 3 3 9 22]
Art Practice: Craft Required Disclosure [ 0 1 3 6 8]
Art Practice: Music Required Disclosure [ 0 1 2 6 7]
Art Practice: Digital Required Disclosure [ 0 2 2 2 8]
Art Practice: Not Artist Required Disclosure [ 0 1 1 5 3]
Art Practice: Other Required Disclosure [ 0 0 3 2 4]
Art Practice: Tattoo Required Disclosure [ 0 0 1 3 4]
Art Practice: Illustration Required Disclosure [ 1 1 2 0 5]
Art Practice: Drawing Required Disclosure [ 0 0 1 4 2]
Art Practice: Sculpture Required Disclosure [ 0 0 0 1 5]
Art Practice Maker Required Disclosure [ 0 0 0 2 2]]
Art Practice: Paint AI Positive for Art [[23 26 33 35 44]
Art Practice: Photography AI Positive for Art [ 8 14 29 35 23]
Art Practice: Writing AI Positive for Art [12 7 12 7 8]
Art Practice: Designer AI Positive for Art [ 5 7 7 11 10]
Art Practice: Craft AI Positive for Art [ 2 4 12 0 0]
Art Practice: Music AI Positive for Art [ 2 3 5 4 2]
Art Practice: Digital AI Positive for Art [ 0 1 10 0 3]
Art Practice: Not Artist AI Positive for Art [ 2 1 2 5 0]
Art Practice: Other AI Positive for Art [ 1 1 4 3 0]
Art Practice: Tattoo AI Positive for Art [ 0 0 3 0 5]
Art Practice: Illustration AI Positive for Art [ 2 1 2 2 2]
Art Practice: Drawing AI Positive for Art [ 0 0 3 2 2]
Art Practice: Sculpture AI Positive for Art [ 1 3 1 0 1]
Art Practice Maker AI Positive for Art [ 2 1 1 0 0]]
Art Practice: Paint AI Threat to Art [[ 9 23 27 54 48]
Art Practice: Photography AI Threat to Art [ 5 17 22 35 30]
Art Practice: Writing AI Threat to Art [ 5 5 8 14 14]
Art Practice: Designer AI Threat to Art [ 4 4 6 15 11]
Art Practice: Craft AI Threat to Art [ 1 3 1 8 5]
Art Practice: Music AI Threat to Art [ 0 2 2 8 4]
Art Practice: Digital AI Threat to Art [ 1 1 6 5 1]
Art Practice: Not Artist AI Threat to Art [ 0 5 1 2 2]
Art Practice: Other AI Threat to Art [ 0 1 2 4 2]
Art Practice: Tattoo AI Threat to Art [ 0 1 1 3 3]
Art Practice: Illustration AI Threat to Art [ 2 1 0 2 4]
Art Practice: Drawing AI Threat to Art [ 1 1 1 3 1]
Art Practice: Sculpture AI Threat to Art [ 0 1 2 1 2]
Art Practice Maker AI Threat to Art [ 0 0 1 1 2]]
Table 21. Contingency table for the secondary variable of interest (Art Practice) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (Categorical) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Compensation: Tax Owner AI User [[ 0 2 2 2 0]
Compensation: None listed Owner AI User [20 9 9 2 2]
Compensation: No profit needed but no profit for companies Owner AI User [26 26 26 20 8]
Compensation: Flat fee Owner AI User [ 5 13 14 8 14]
Compensation: Donate art to train models Owner AI User [ 4 6 12 26 33]
Compensation: No profit needed but no profit for anyone Owner AI User [ 5 9 3 7 3]
Compensation: Other Owner AI User [ 3 0 4 1 2]
Compensation: Portion from users derivatives Owner AI User [ 3 13 5 8 3]
Compensation: Portion of profit from anyone Owner AI User [ 4 7 5 9 7]
Compensation: Portion of profit from AI company Owner AI User [16 14 9 7 5]
Compensation: Don’t need profit Owner AI User [ 1 2 6 6 7]]
Compensation: Tax Owner Artist [[ 0 1 1 3 1]
Compensation: None listed Owner Artist [ 6 10 10 5 11]
Compensation: No profit needed but no profit for companies Owner Artist [11 32 29 23 11]
Compensation: Flat fee Owner Artist [ 5 16 13 11 9]
Compensation: Donate art to train models Owner Artist [13 17 8 17 26]
Compensation: No profit needed but no profit for anyone Owner Artist [ 1 9 3 9 5]
Compensation: Other Owner Artist [ 1 3 4 1 1]
Compensation Portion from users derivatives Owner Artist [ 0 9 11 9 3]
Compensation: Portion of profit from anyone Owner Artist [ 1 8 10 6 7]
Compensation: Portion of profit from AI company Owner Artist [ 1 9 18 12 10]
Compensation Owner Artist [ 4 6 5 4 3]]
Compensation: Tax Owner AI Company [[ 0 3 1 2 0]
Compensation: None listed Owner AI Company [20 10 8 2 2]
Compensation: No profit needed but no profit for companies Owner AI Company [39 33 18 10 6]
Compensation: Flat fee Owner AI Company [10 12 12 13 7]
Compensation: Donate art to train models Owner AI Company [18 9 12 18 24]
Compensation: No profit needed but no profit for anyone Owner AI Company [ 9 10 2 4 2]
Compensation: Other Owner AI Company [ 6 1 3 0 0]
Compensation Portion from users derivatives Owner AI Company [11 11 3 4 3]
Compensation: Portion of profit from anyone Owner AI Company [ 6 8 10 5 3]
Compensation: Portion of profit from AI company Owner AI Company [23 12 6 7 3]
Compensation: Portion of profit from AI company Owner AI Company [ 4 4 7 4 3]]
Table 22. Contingency table for the secondary variable of interest (Fair compensation) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (Categorical) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Compensation: Tax Required Disclosure [[ 0 0 0 4 2]
Compensation: None listed Required Disclosure [ 1 0 7 7 27]
Compensation: No profit needed but no profit for companies Required Disclosure [ 2 4 13 28 56]
Compensation: Flat fee Required Disclosure [ 2 4 2 26 19]
Compensation: Donate art to train models Required Disclosure [ 3 8 8 27 35]
Compensation: No profit needed but no profit for anyone Required Disclosure [ 0 2 1 10 14]
Compensation: Other Required Disclosure [ 0 2 4 2 2]
Compensation Portion from users derivatives Required Disclosure [ 0 1 1 13 16]
Compensation: Portion of profit from anyone Required Disclosure [ 2 1 6 9 14]
Compensation: Portion of profit from AI company Required Disclosure [ 0 0 5 17 29]
Compensation: Portion of profit from AI company Required Disclosure [ 3 2 8 3 6]]
Compensation: Tax AI Positive for Art [[ 0 1 2 1 2]
Compensation: None listed AI Positive for Art [17 8 12 2 3]
Compensation: No profit needed but no profit for companies AI Positive for Art [14 21 34 17 17]
Compensation: Flat fee AI Positive for Art [ 3 9 20 13 8]
Compensation: Donate art to train models AI Positive for Art [ 1 1 7 29 42]
Compensation: No profit needed but no profit for anyone AI Positive for Art [ 6 2 6 9 4]
Compensation: Other AI Positive for Art [ 3 0 5 0 2]
Compensation Portion from users derivatives AI Positive for Art [ 3 8 8 9 3]
Compensation: Portion of profit from anyone AI Positive for Art [ 4 7 8 7 6]
Compensation: Portion of profit from AI company AI Positive for Art [ 8 10 16 11 6]
Compensation: Portion of profit from AI company AI Positive for Art [ 1 2 6 6 7]]
Compensation: Tax AI Threat to Art [[ 0 1 4 1 0]
Compensation: None listed AI Threat to Art [ 2 2 7 11 20]
Compensation: No profit needed but no profit for companies AI Threat to Art [ 3 16 15 43 26]
Compensation: Flat fee AI Threat to Art [ 5 8 6 22 12]
Compensation: Donate art to train models AI Threat to Art [10 13 14 20 23]
Compensation: No profit needed but no profit for anyone AI Threat to Art [ 2 5 4 12 4]
Compensation: Other AI Threat to Art [ 0 2 5 2 1]
Compensation: Portion from users derivatives AI Threat to Art [ 0 5 3 13 10]
Compensation: Portion of profit from anyone AI Threat to Art [ 3 5 5 8 11]
Compensation: Portion of profit from AI company AI Threat to Art [ 0 7 9 18 17]
Compensation: Portion of profit from AI company AI Threat to Art [ 3 1 8 5 5]]
Table 23. Contingency table for the secondary variable of interest (Fair compensation) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Secondary Variable (Categorical) Dependent Variable (Likert) Contingency Table [SD D N A SA]
Gender Male Owner AI User [[36 54 61 62 60]
Gender Female Owner AI User [47 39 29 25 21]
Gender Non-binary Owner AI User [ 4 5 4 7 3]]
Gender Male Owner Artist [[23 66 63 63 57]
Gender Female Owner Artist [18 43 40 32 28]
Gender Non-binary Owner Artist [ 2 7 7 5 2]]
Gender Male Owner AI Company [[71 59 60 43 40]
Gender Female Owner AI Company [65 41 20 23 12]
Gender Non-binary Owner AI Company [ 9 8 2 3 1]]
Gender Male Required Disclosure [[ 6 15 35 95 122]
Gender Female Required Disclosure [ 6 7 17 45 86]
Gender Non-binary Required Disclosure [ 1 2 2 6 12]]
Gender Male AI Positive for Art [[21 41 80 68 62]
Gender Female AI Positive for Art [35 26 38 29 33]
Gender Non-binary AI Positive for Art [ 4 2 6 6 5]]
Gender Male AI Threat to Art [[13 42 49 92 76]
Gender Female AI Threat to Art [12 20 28 51 50]
Gender Non-binary AI Threat to Art [ 3 3 2 12 3]]
Table 24. Contingency table for the secondary variable of interest (Gender) and dependent variables on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Strongly Disagree Disagree Neutral Agree Strongly Agree
Strongly Disagree 1 0 3 2 22
Disagree 3 4 9 29 20
 Neutral 3 6 41 17 12
Agree 11 29 59 37 18
Strongly Disagree 42 29 12 18 27
Table 25. Vertical side of table represents AI as a threat to the art workforce and the horizontal side of the table represents AI as a positive development for the field of art. Contingency table for the AI positive development for art and AI threat to art workforce on a 5-point Likert scale strongly disagree [(SD, leftmost number in bracket), disagree (D), neutral (N), agree (A), and strongly agree (SA, rightmost number in bracket)]. Outer brackets indicate the start and end of each contingency table, and the inner brackets indicate the start and end of a row in the table.
Appendix F Tests for Independence
F.0.1. Kruskal-Wallis test for independence:
To test for independence between dependent variables that were ordinal with variables of interest that are categorical, we ran Kruskall-Wallis tests for independence. The results of these tests can be seen in Tables 26, 27, 30, 31, 28, 29.
Variable p-value H Statistic Degrees of Freedom
Artist 0.9873871802 0.0002499081613 1
Gender 0.01738441229 8.104362639 2
Art Practice 0.007501359367 28.5814306 13
AI Model Familiarity 0.9798888491 0.4306800685 4
Age 0.03938656365 8.345451383 3
Compensation 1.05E-16 98.58439653 10
Table 26. Results of the Kruskall-Wallis test for independence for the dependent variable (5-point Likert) AI art models are a positive development in the field of art. Strong evidence follows a threshold of a p-value less than 0.05.
Variable p-value H Statistic Degrees of Freedom
Artist 0.5445539432 0.3671644759 1
Gender 0.5726378558 1.115003553 2
Art Practice 0.9024842342 6.993538631 13
AI Model Familiarity 0.4531083849 3.665876796 4
Age 0.1005912351 6.237926759 3
Compensation 0.04021463347 19.003785 10
Table 27. Results of the Kruskall-Wallis test for independence for the dependent variable (5-point Likert) AI models are a threat to art workers. Strong evidence follows a threshold of a p-value less than 0.05.
Variable p-value H Statistic Degrees of Freedom
Artist 0.8660551636 0.02845002072 1
Gender 0.326366829 2.239466576 2
Art Practice 0.8988829186 7.062863624 13
AI Model Familiarity 7.67E-05 24.08859461 4
Age 0.1962046718 4.686999195 3
Compensation 0.001618196457 28.29622161 10
Table 28. Results of the Kruskall-Wallis test for independence for the dependent variable (5-point Likert) model creators should be required to disclose in detail what art & images they use to train their AI models. Strong evidence follows a threshold of a p-value less than 0.05.
Variable p-value H Statistic Degrees of Freedom
Artist 0.795701369 0.06703661173 1
Gender 5.41E-05 19.65071632 2
Art Practice 0.1586078361 17.97141505 13
AI Model Familiarity 0.1628188729 6.531460606 4
Age 0.4668168207 2.547225387 3
Compensation 5.36E-14 84.94217137 10
Table 29. Results of the Kruskall-Wallis test for independence for the dependent variable (5 point Likert), if an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) that work and its derivatives, should be considered the property of the person who used the AI model to generate the artwork. Strong evidence follows a threshold of a p-value less than 0.05.
Variable p-value H Statistic Degrees of Freedom
Artist 0.2607376687 1.264836076 1
Gender 0.2515206604 2.760460289 2
Art Practice 0.05775284946 21.84857601 13
AI Model Familiarity 0.2406398118 5.489660662 4
Age 0.261387791 4.000707054 3
Compensation 0.3765939846 10.75907326 10
Table 30. Results of the Kruskall-Wallis test for independence for the dependent variable (5 point Likert) If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh) that work and its derivatives should be considered the property of the artist. Strong evidence follows a threshold of a p-value less than 0.05.
Variable p-value H Statistic Degrees of Freedom
Artist 0.3626709363 0.8286232383 1
Gender 0.0005455491796 15.027435 2 2
Art Practice 0.0004109645224 37.0224258 13
AI Model Familiarity 4.31E-06 30.27078455 4
Age 0.4650478321 2.557169371 3
Compensation 1.34E-07 51.62344111 10
Table 31. Results of the Kruskall-Wallis test for independence for the dependent variable (5 point Likert), if an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) that work and its derivatives, should be considered the property of the AI model creators. Strong evidence follows a threshold of a p-value less than 0.05.
F.0.2. Cochran-Armitage test for independence:
To test for independence between dependent variables that were ordinal with variables of interest that are binary, we ran Cochran-Armitage tests for independence. The results of these tests can be seen in Tables 32, 33, 36, 37, 34, and 35.
Variable p-value Statistic Null mean Null SD Z-score
Person of color 0.00206563 243 208.5620609 11.17886095 3.08063042
Professional Artist 0.08792146 472 448.8618267 13.55912215 1.706465436
Purchased Art 0.82483924 855 856.9180328 8.666129548 -0.2213251921
Used AI art 0.00000000 603 489.6674473 13.59406494 8.336914175
Table 32. Results of the Cochran-Armitage test for independence for the dependent variable (5-point Likert) AI art models are a positive development in the field of art. Strong evidence follows a threshold of a p-value less than 0.05.
Variable p-value Statistic Null mean Null SD Z-score
Person of color 0.27332660 254 242.6042155 10.40298712 1.09543388
Professional Artist 0.33653136 510 522.1264637 12.61804523 -0.9610413879
Purchased Art 0.47302764 991 996.7868852 8.064652959 -0.7175615956
Used AI art 0.00031337 524 569.5925059 12.6505628 -3.60399032
Table 33. Results of the Cochran-Armitage test for independence for the dependent variable (5-point Likert) AI models are a threat to art workers. Strong evidence follows a threshold of a p-value less than 0.05.
Variable p-value Statistic Null mean Null SD Z-score
Person of color 0.97821917 293 293.236534 8.663735458 -0.0273016136
Professional Artist 0.08842447 649 631.0960187 10.50846306 1.703767826
Purchased Art 0.22323397 1213 1204.819672 6.716342048 1.217973684
Used AI art 0.06462021 669 688.4683841 10.53554409 -1.847876475
Table 34. Results of the Cochran-Armitage test for independence for the dependent variable (5-point Likert) model creators should be required to disclose in detail what art & images they use to train their AI models. Strong evidence follows a threshold of a p-value less than 0.05.
Variable p-value Statistic Null mean Null SD Z-score
Person of color 0.00746442 214 182.2763466 11.85761676 2.675381911
Professional Artist 0.98389084 392 392.2903981 14.38240218 -0.02019121164
Purchased Art 0.15992987 736 748.9180328 9.192317846 -1.405307454
Used AI art 0.00001229 491 427.9531616 14.41946662 4.372341923
Table 35. Results of the Cochran-Armitage test for independence for the dependent variable (5 point Likert), if an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) that work and its derivatives, should be considered the property of the person who used the AI model to generate the artwork. Strong evidence follows a threshold of a p-value less than 0.05.
Variable p-value Statistic Null mean Null SD Z-score
Person of color 0.00037047 237 198.4355972 10.83186977 3.560272014
Professional Artist 0.59775878 434 427.0679157 13.13824781 0.5276262413
Purchased Art 0.74883816 818 815.3114754 8.397133403 0.3201717135
Used AI art 0.08222217 443 465.8922717 13.17210597 -1.737935582
Table 36. Results of the Cochran-Armitage test for independence for the dependent variable (5 point Likert) If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh) that work and its derivatives should be considered the property of the artist. Strong evidence follows a threshold of a p-value less than 0.05.
Variable p-value Statistic Null mean Null SD Z-score
Person of color 0.00001507 188 137.030445 11.77770463 4.327630607
Professional Artist 0.02548492 263 294.9133489 14.28547474 -2.233971885
Purchased Art 0.18814356 551 563.0163934 9.130367983 -1.316090815
Used AI art 0.68942716 316 321.7236534 14.32228939 -0.3996325755
Table 37. Results of the Cochran-Armitage test for independence for the dependent variable (5 point Likert), if an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) that work and its derivatives, should be considered the property of the AI model creators. Strong evidence follows a threshold of a p-value less than 0.05.